# When LLMs Meets Acoustic Landmarks: An Efficient Approach to Integrate Speech into Large Language Models for Depression Detection

### v1

@@ Abstract
Depression is a critical concern in global mental health, prompting extensive research into AI-based detection methods. Among various AI technologies, Large Language Models (LLMs) stand out for their versatility in mental healthcare applications. However, their primary limitation arises from their exclusive dependence on textual input, which constrains their overall capabilities. Furthermore, the utilization of LLMs in identifying and analyzing depressive states is still relatively untapped. In this paper, we present an innovative approach to integrating acoustic speech information into the LLMs framework for multimodal depression detection. We investigate an efficient method for depression detection by integrating speech signals into LLMs utilizing Acoustic Landmarks. By incorporating acoustic landmarks, which are specific to the pronunciation of spoken words, our method adds critical dimensions to text transcripts. This integration also provides insights into the unique speech patterns of individuals, revealing the potential mental states of individuals. Evaluations of the proposed approach on the DAIC-WOZ dataset reveal state-of-the-art results when compared with existing Audio-Text baselines. In addition, this approach is not only valuable for the detection of depression but also represents a new perspective in enhancing the ability of LLMs to comprehend and process speech signals. 111code will be made public upon publication
@@ 2.1 Large Language Models
Large language models have achieved success in natural language processing and have been extended to encompass computer vision and speech signal processing Brown et al. (2020); Touvron et al. (2023). However, there is a significant gap in research aimed at enabling LLMs to comprehend speech efficiently.
@@ 2.2 Acoustic Landmarks
The concept of acoustic landmarks originally stems from research on distinctive features Garvin (1953). Some researchers posit that for certain phonetic contrasts, a listener relies on acoustic landmarks to gather the necessary acoustic cues for deciphering the underlying distinctive features Liu (1996). This perspective highlights the importance of these landmarks in the auditory processing and interpretation of speech. Subsequent research has utilized acoustic landmarks for applications in speech recognition Liu (1996); He et al. (2019) as well as in addressing mental health-related problems Huang et al. (2018, 2019a). Although different scholars have slightly varied definitions of acoustic landmarks, Joel and colleagues Boyce et al. (2012) expanded upon Liu’s paper Liu (1996) by releasing a MATLAB version of a landmark detection toolkit, which has become the most widely used version of landmark technology.
@@ 2.3 Automatic Depression Detection
The use of AI technology for depression detection has been developing for many years. Some researchers  Cummins et al. (2011); Huang et al. (2018, 2019a) have utilized traditional methods such as Support Vector Machines (SVMs) Noble (2006) for depression detection. With the advancement of deep learning technologies, an increasing number of researchers have been experimenting with deep learning approaches for depression detection. Zhao and others have explored the use of transformer models for processing speech inputs in depression detection Zhao et al. (2020). Shen and colleagues have employed BI-LSTM architectures, combining text and speech for this purpose Shen et al. (2022). Further extending these techniques, Wu Wu et al. (2023) utilized speech self-supervised models Chen et al. (2022); Hsu et al. (2021) and integrated them with RoBERTa Liu et al. (2019) for a more comprehensive text-audio multimodal approach to depression detection.
@@ 3.2.1 Landmarks Extraction
g
vibration of vocal folds start (+) or end (–)
b
onset (+) or offset (–) of existence of turbulent noise during obstruent regions
s
releases (+) or closures (–) of a nasal
v
voiced frication onset (+) or offset (–)
p
periodicity start (+) or end (–)
f
frication onset (+) or offset (–)
@@ 4.2 Main Result: Performance of different LLMs in Depression Detection task
The introduction of landmarks led to enhanced performance across all models, affirming the effectiveness of our method in integrating landmarks. Landmarks can represent some of the acoustic information due to affective variation, providing additional information that assists LLMs in detecting depression. Nonetheless, the efficacy of using landmarks in isolation for depression detection was found to be suboptimal. Drawing on past research, we believe this is due to the fact that even after cross-modal instruction fine-tuning, relying solely on information from other modalities (such as audio or visual) could potentially impair the stability of LLMs Zhang et al. (2023); Li et al. (2023). When we combined multiple Llama2 models that had integrated both text and landmark information for depression detection, we achieved SOTA results as shown in table 3. Furthermore, as indicated in Table 3, there is a gradual improvement in Llama2’s performance in depression detection tasks as the number of sub-dialogues per dialogue increases. This observation further emphasizes the crucial role that data quantity plays in the effectiveness of depression detection tasks.
@@ 5.1 Effect of Hint in Cross-Modal Instruction Fine-Tuning
During the Cross-Modal Instruction Fine-Tuning phase, we discovered that providing a hint to the LLMs is crucial. In other words, informing the LLMs whether the data sample originates from a patient with depression significantly impacts the training outcome. As evident from Figure 4, without a hint, the loss converged to around 1.76 (as shown in Figure 3(a)). In contrast, with a hint, the loss consistently converged to near 1.1 (as depicted in Figures 3(b) and 3(c)). Figure 3(d) offers a more vivid illustration of the substantial difference that the presence or absence of a hint makes to the model’s performance in our empirical study. This phenomenon supports our previous conjecture that individuals with depression and those who are healthy differ in their vocal expressions and that landmarks are capable of reflecting this characteristic. Although the differences between Llama2 and Llama2 Chat are not substantial, it is still observable that, in this phase, Llama2 outperforms its Chat version. We will provide a more detailed discussion in the subsequent section.
@@ 5.2 How LLMs Learn from Acoustic Landmarks
To further investigate how LLMs learn acoustic landmarks, we extended the application of LoRA beyond just the attention layers, applying it across all layers for comprehensive analysis Pu et al. (2023); Sun et al. (2023). To find the matrix with the greatest contribution, we first need to define the method for calculating the contribution of a matrix. We can approximately consider the changes in the LoRA matrix as indicative of its contribution to the task He et al. (2021). Therefore, we assess that the contribution of a matrix is calculated by summing the absolute values of all its elements, normalized by the total number of elements in the matrix. Suppose we have a set of LoRA matrices $L_{1},L_{2},\ldots,L_{n}$, each matrix $L_{i}$ being an $a\times b$ matrix. Then, the contribution $C_{i}$ of matrix $L_{i}$ can be calculated using the formula:
@@ Acknowledgement
This work was supported by Australian Research Council Discovery Project DP230101184. Additionally, we extend our gratitude to Hanyu Meng for assistance in creating the graphs for this paper.
@@ A.1 General Processing Details
$L^{(cp)}_{b}[n]=10\cdot\log_{10}\left(\frac{1}{2N_{cp}+1}\sum_{k=-N_{cp}}^{N_{cp}}E_{b}[n+k]\right),$ (4)
@@
$L^{(fp)}_{b}[n]=10\cdot\log_{10}\left(\frac{1}{2N_{fp}+1}\sum_{k=-N_{fp}}^{N_{fp}}E_{b}[n+k]\right),$ (7)
@@ Appendix C Sample of Hint Cross-modal Instruction Fine Tuning
corresponding to these transcripts
@@
Please try to predict the concatenated acoustic landmarks corresponding to these transcripts.
@@ Appendix D Sample of Instruction Fine-Tuning for Depression Detection
 "Categorize these dialogues as either depression or healthy based on the its transcripts.
@@
 "Categorize these dialogues as either depression or healthy based on the its acoustic landmarks.
@@
"Categorize these dialogues as either depression
or healthy based on the its transcripts and acoustic landmarks.
@@
### Acoustic Landmark:{landmark}

### v2

# When LLMs Meet Acoustic Landmarks: An Efficient Approach to Integrate Speech into Large Language Models for Depression Detection

@@ Abstract
Depression is a critical concern in global mental health, prompting extensive research into AI-based detection methods. Among various AI technologies, Large Language Models (LLMs) stand out for their versatility in mental healthcare applications. However, their primary limitation arises from their exclusive dependence on textual input, which constrains their overall capabilities. Furthermore, the utilization of LLMs in identifying and analyzing depressive states is still relatively untapped. In this paper, we present an innovative approach to integrating acoustic speech information into the LLMs framework for multimodal depression detection. We investigate an efficient method for depression detection by integrating speech signals into LLMs utilizing Acoustic Landmarks. By incorporating acoustic landmarks, which are specific to the pronunciation of spoken words, our method adds critical dimensions to text transcripts. This integration also provides insights into the unique speech patterns of individuals, revealing the potential mental states of individuals. Evaluations of the proposed approach on the DAIC-WOZ dataset reveal state-of-the-art results when compared with existing Audio-Text baselines. In addition, this approach is not only valuable for the detection of depression but also represents a new perspective in enhancing the ability of LLMs to comprehend and process speech signals.
@@ 2.1 Large Language Models
Large language models have achieved success in natural language processing and have been extended to encompass computer vision and speech signal processing Brown et al. (2020); Touvron et al. (2023); Li et al. (2023b); Liu et al. (2024). However, there is a significant gap in research aimed at enabling LLMs to comprehend speech efficiently.
@@ 2.2 Acoustic Landmarks
The concept of acoustic landmarks originally stems from research on distinctive features Garvin (1953); Zhang et al. (2024a). Some researchers posit that for certain phonetic contrasts, a listener relies on acoustic landmarks to gather the necessary acoustic cues for deciphering the underlying distinctive features Liu (1996). This perspective highlights the importance of these landmarks in the auditory processing and interpretation of speech. Subsequent research has utilized acoustic landmarks for applications in speech recognition Liu (1996); He et al. (2019) as well as in addressing mental health-related problems Huang et al. (2018, 2019a). Although different scholars have slightly varied definitions of acoustic landmarks, Joel and colleagues Boyce et al. (2012) expanded upon Liu’s paper Liu (1996) by releasing a MATLAB version of a landmark detection toolkit, which has become the most widely used version of landmark technology.
@@ 2.3 Automatic Depression Detection
The use of AI technology for depression detection has been developing for many years. Some researchers  Cummins et al. (2011); Huang et al. (2018, 2019a) have utilized traditional methods such as Support Vector Machines (SVMs) Noble (2006) for depression detection. With the advancement of deep learning technologies Gulati et al. (2020); Zhang et al. (2024c), an increasing number of researchers have been experimenting with deep learning approaches for depression detection. Zhao and others have explored the use of transformer models for processing speech inputs in depression detection Zhao et al. (2020). Shen and colleagues have employed BI-LSTM architectures, combining text and speech for this purpose Shen et al. (2022). Further extending these techniques, Wu Wu et al. (2023) utilized speech self-supervised models Chen et al. (2022); Hsu et al. (2021); Liu et al. (2022) and integrated them with RoBERTa Liu et al. (2019) for a more comprehensive text-audio multimodal approach to depression detection.
@@ 3.2.1 Landmarks Extraction
g vibration of vocal folds start (+) or end (–)
b onset (+) or offset (–) of existence of turbulent noise during obstruent regions
s releases (+) or closures (–) of a nasal
v voiced frication onset (+) or offset (–)
p periodicity start (+) or end (–)
f frication onset (+) or offset (–)
@@ 4.2 Main Result: Performance of different LLMs in Depression Detection task
The introduction of landmarks led to enhanced performance across all models, affirming the effectiveness of our method in integrating landmarks. Landmarks can represent some of the acoustic information due to affective variation, providing additional information that assists LLMs in detecting depression. Nonetheless, the efficacy of using landmarks in isolation for depression detection was found to be suboptimal. Drawing on past research, we believe this is due to the fact that even after cross-modal instruction fine-tuning, relying solely on information from other modalities (such as audio or visual) could potentially impair the stability of LLMs Zhang et al. (2023); Li et al. (2023c). When we combined multiple Llama2 models that had integrated both text and landmark information for depression detection, we achieved SOTA results as shown in table 3. Furthermore, as indicated in Table 3, there is a gradual improvement in Llama2’s performance in depression detection tasks as the number of sub-dialogues per dialogue increases. This observation further emphasizes the crucial role that data quantity plays in the effectiveness of depression detection tasks.
@@ 5.1 Effect of Hint in Cross-Modal Instruction Fine-Tuning
During the Cross-Modal Instruction Fine-Tuning phase, we discovered that providing a hint to the LLMs is crucial. In other words, informing the LLMs whether the data sample originates from a patient with depression significantly impacts the training outcome. As evident from Figure 4, without a hint, the loss converged to around 1.76 (as shown in Figure 4(a)). In contrast, with a hint, the loss consistently converged to near 1.1 (as depicted in Figures 4(b) and 4(c)). Figure 4(d) offers a more vivid illustration of the substantial difference that the presence or absence of a hint makes to the model’s performance in our empirical study. This phenomenon supports our previous conjecture that individuals with depression and those who are healthy differ in their vocal expressions and that landmarks are capable of reflecting this characteristic. Although the differences between Llama2 and Llama2 Chat are not substantial, it is still observable that, in this phase, Llama2 outperforms its Chat version. We will provide a more detailed discussion in the subsequent section.
@@ 5.2 How LLMs Learn from Acoustic Landmarks
To further investigate how LLMs learn acoustic landmarks, we extended the application of LoRA beyond just the attention layers, applying it across all layers for comprehensive analysis Pu et al. (2023); Sun et al. (2023); Li et al. (2023a); Zhang et al. (2024b). To find the matrix with the greatest contribution, we first need to define the method for calculating the contribution of a matrix. We can approximately consider the changes in the LoRA matrix as indicative of its contribution to the task He et al. (2021). Therefore, we assess that the contribution of a matrix is calculated by summing the absolute values of all its elements, normalized by the total number of elements in the matrix. Suppose we have a set of LoRA matrices $L_{1},L_{2},\ldots,L_{n}$, each matrix $L_{i}$ being an $a\times b$ matrix. Then, the contribution $C_{i}$ of matrix $L_{i}$ can be calculated using the formula:
@@ Acknowledgement
This work was supported by Australian Research Council Discovery Project DP230101184.
@@ A.1 General Processing Details
$L^{(cp)}_{b}[n]=10\cdot\log_{10}\left(\frac{1}{2N_{cp}+1}\sum_{k=-N_{cp}}^{N_{cp}}E_{b}[n+k]\right)$
@@
$L^{(fp)}_{b}[n]=10\cdot\log_{10}\left(\frac{1}{2N_{fp}+1}\sum_{k=-N_{fp}}^{N_{fp}}E_{b}[n+k]\right)$
@@ Appendix C Sample of Hint Cross-modal Instruction Fine Tuning
corresponding to these transcripts.
@@
Please try to predict the concatenated acoustic landmarks
corresponding to these transcripts.
@@ Appendix D Sample of Instruction Fine-Tuning for Depression Detection
"Categorize these dialogues as either depression or healthy based on its transcripts.
@@
"Categorize these dialogues as either depression or healthy based on its acoustic landmarks.
@@
"Categorize these dialogues as either depression or healthy based on its transcripts and acoustic landmarks.
@@
### Acoustic Landmark:{landmarks}